OpenStudy (anonymous):

Solve for x: 3 over 3 x plus 1 over quantity x plus 4 equals 10 over 7x

OpenStudy (anonymous):

\[\frac{ x }{ 3x } + \frac{ 1 }{ x + 4 } = \frac{ 10 }{7x }\]

OpenStudy (anonymous):

Answer choices x = –3 x = 3 x = –3 and x = 3 No Solution

OpenStudy (anonymous):

first start by finding a common denominator for the left side of the equation can u do that?

OpenStudy (anonymous):